# SADA: Semantic adversarial unsupervised domain adaptation for Temporal Action Localization

### 20 Dec 2023 v1

@@ F Main ablations on other scenarios
In this section, we extend the ablation studies presented in Sec. 4.3 by analyzing two additional scenarios. Concretely, we compare the performance of SADA with the adaptations of different UDA methods, and with different variations of our loss, in Scenario 2 and 4.
F.1 Domain adaptation methods
In Tab. B we present the comparison of our proposed method to the other considered UDA baselines. This indicates that the first scenario (top), namely white new videos, poses important challenges when aligning its corresponding source and target domain. More specifically, in this scenario, all the considered UDA baselines worsen the performance of the source-only variant. The behavior of the baselines in the second scenario (down), namely white old, is similar. In this case, however, DANN [17] does improve the source-only variant. This is consistent with our observations from Sec. 4.2 where DANN [17] was also the best-performing baseline. Similarly, SSTDA [8] is the worst performing method in both scenarios, especially degrading in the second scenario (bottom) by up to a relative . In contrast, our method improves all the baselines in both scenarios, attaining up to a of relative improvement in the second scenario.
@@ F.1 Domain adaptation methods
| Semantic centroids [58] | 30.17 | 27.32 | 18.55 | 25.34 |
| SSTDA [8] | 23.45 | 20.95 | 13.37 | 19.26 |
@@
F.2 Ablation over variants of the SADA loss
@@
Similarly to Sec. 4.3, in Tab. C we ablate over different variants of our loss on Scenario 2 and 4. These scenarios define the source domain as white old and new counter kitchens and the rest as the target, respectively. Observe in this table that our main insights from Sec. 4.3 are mostly confirmed. Concretely, one of our main observations was that aligning the complete global distribution was generally preferable over aligning the partial background distribution only. This observation is satisfied in most of the scenarios that we study –with the exception of Scenario 3. Similarly, in Sec. 4.3 we argued that combining alignment loss that presented overlaps yielded also a decrease in performance. More in detail, this happens when we couple either the local or the background loss with a global alignment approach. In these cases, the embeddings have a double (simultaneous) alignment objective, which we argue is not desirable. As observed in Tab. C, this observation is also mostly satisfied in these scenarios as the combination of these losses yields a performance degradation in 7 of the 8 cases studied in this paper – included in Tab. C and Tab. 4 of the main paper.
G Ablation per class
| take | put | wash | open | close | insert | turn-on | cut | turn-off | pour | |
| Src-only | 24.20 | 29.30 | 37.30 | 32.69 | 23.89 | 10.98 | 37.39 | 18.18 | 26.24 | 18.55 |
| SADA (ours) | 26.31 | 30.70 | 39.14 | 34.68 | 23.42 | 10.58 | 39.86 | 17.05 | 26.58 | 18.78 |
| take | put | wash | open | close | insert | turn-on | cut | turn-off | pour | |
| Src-only | 24.11 | 25.69 | 29.77 | 31.96 | 27.8 | 6.69 | 25.52 | 34.16 | 17.48 | 29.18 |
| SADA (ours) | 28.32 | 31.25 | 31.18 | 34.52 | 31.82 | 7.29 | 26.66 | 39.12 | 18.68 | 28.34 |
In this section, we complement the analysis from Sec. 4.2 with the detailed class-wise metrics. Concretely, in Tab. D and Tab. E we show the respective class-wise mAP scores of the 10 considered classes on both Scenario 1 and 3. For the analysis, we also include the results obtained with the source-only variant of our model as well as our chosen DANN[17] baseline.
Observe that in Scenario 3 (see Tab. D) our model attains the best class-wise performance in 5 of the classes, while DANN [17] does so on 3, and the source-only model on 2. In contrast, in Scenario 1 (see Tab. E) our method obtains a much clearer improvement over the chosen baselines, yielding the best results in 8 of the 10 classes. Overall, we can observe that our method performs very well in the 3 majority classes of both scenarios, where we highlight the relative improvement of our model over DANN [17] of for the class put in Scenario 3. Moreover, our method fails to improve the performance of action pour in both scenarios, which we attribute to the lack of sufficient data to operate on our proposed methodology. We also highlight that as observed in Sec.4.2, Scenario 1 presents a more challenging setup, degrading the performance in other actions such as open, close, insert or cut. This indicates the existence of intrinsic qualitative aspects that harden the adaptation when dealing with old videos. This is not the case in Scenario 3, as aside from the aforementioned pour segments, it only fails to achieve the best results for the turn-on action.
H Ablation of the effect of negative anchors
[…]
| Filtered % | mAP {10,20,30,40,50}% | Avg | ||||
| 0% | 32.69 | 31.49 | 29.17 | 25.51 | 19.72 | 27.72 |
| 25% | 32.54 | 31.87 | 30.06 | 26.99 | 21.91 | 28.67 |
| 50% | 34.46 | 33.55 | 31.29 | 27.59 | 21.18 | 29.61 |
| 75% | 35.09 | 34.53 | 32.97 | 30.05 | 23.97 | 31.32 |
| 100% | 35.57 | 35.24 | 34.16 | 30.72 | 24.50 | 32.04 |
[…]
| Filtered % | mAP {10,20,30,40,50}% | Avg | ||||
| 0% | 31.60 | 30.29 | 28.22 | 24.47 | 18.98 | 26.71 |
| 25% | 32.76 | 31.60 | 29.81 | 26.14 | 20.24 | 28.11 |
| 50% | 35.25 | 34.25 | 32.68 | 29.25 | 23.01 | 30.89 |
| 75% | 38.39 | 37.77 | 36.40 | 33.19 | 26.33 | 34.42 |
| 100% | 43.93 | 43.59 | 42.61 | 39.26 | 31.85 | 40.25 |
[…]
| Filtered % | mAP {10,20,30,40,50}% | Avg | ||||
| 0% | 32.69 | 31.49 | 29.17 | 25.51 | 19.72 | 27.72 |
| 25% | 31.48 | 30.34 | 27.97 | 24.07 | 18.93 | 26.56 |
| 50% | 32.11 | 31.11 | 29.25 | 25.85 | 20.72 | 27.81 |
| 75% | 35.99 | 35.28 | 33.48 | 29.90 | 23.64 | 31.66 |
| 100% | 37.53 | 37.06 | 35.73 | 32.34 | 26.45 | 33.82 |
[…]
| Filtered % | mAP {10,20,30,40,50}% | Avg | ||||
| 0% | 31.60 | 30.29 | 28.22 | 24.47 | 18.98 | 26.71 |
| 25% | 32.25 | 30.46 | 28.04 | 24.54 | 19.03 | 26.86 |
| 50% | 33.43 | 32.06 | 29.98 | 26.44 | 20.55 | 28.49 |
| 75% | 38.26 | 37.33 | 35.72 | 32.53 | 25.61 | 33.89 |
| 100% | 47.08 | 46.69 | 45.65 | 42.31 | 34.42 | 43.23 |
We argue that one of the main challenges that our method faces is how to deal with the excess of anchors predicted as class 0, which as described in Sec. 3.4 we refer to as negative anchors. These are the anchors that are not directly matched with a GT segment. In these cases, hence, we hope for the model to learn to identify these anchor embeddings as background, not predicting any action class during inference. This, however, is a cumbersome issue that most of the anchor-based methods face, which is exacerbated when doing domain adaptation in this complicated setup. In this case, it is not clear the amount of noise that these embeddings induce in the alignment strategy. Even though we leave a more thorough analysis of this issue as future work, in this section we aim to provide minimal proof of the potential negative effect that this has on the model’s performance.
Concretely, we propose an experiment where we progressively mask out the set of negative anchors. For this, in Tab. F we randomly mask out these only during inference. In other words, we leave all the training routines unchanged – thus considering all the anchors – but ensure that a percentage of the negative anchors are not considered during inference. This measures the predicting potential of a model that is able to perfectly deal with these embeddings. As observed in Tab. F, in both considered scenarios we observe a linear performance increase, reaching up to a relative improvement of the version with filtered negative anchors over the version. This is, the variant where only the positive anchors are considered over that containing both positive and all the negative ones.
In Tab. G we propose an alternative analysis. In this case, we do not only mask the negative anchors during inference but also during training. In this case, for instance, masking out of the negative anchors additionally implies not training any of the task losses on them, nor using them on our alignment loss. As observed in the results, masking them out entirely – i.e., – has an even bigger impact, reaching a relative improvement of over the version. Nevertheless, in this case, we observe that the improvement is no longer linear, and even worsens the performance for small percentages of masking. We hypothesize this is induced by a reduction of the training data of background class. In other words, while masking some negative embeddings has a positive effect, we must also consider the negative effect that having less training data has on the ones that we do consider. Hence, until this trade-off is dominated by the positive effect of masking out the majority of negative anchors, the improvement remains little to non-existent. In conclusion, these experiments emphasize the need for devising carefully designed models that can properly cope with this intrinsic limitations of anchor-based methods.

### 05 Apr 2024 v2

@@ Appendix D Extended results and ablations
In this section, we extend the ablation studies presented in Sec. IV.C of the main text by analyzing two additional scenarios. Concretely, we compare the performance of SADA with the adaptations of different UDA methods, and with different variations of our loss, when evaluated in scenarios S2 and S4.
D.1 Domain adaptation methods
In Table S5 we present the comparison of our proposed method to the other considered UDA baselines. Concretely, the first scenario (left) refers to the white new videos, while the second scenario (right) refers to the white old videos. Similarly to Sec. IV.C of the main text, we observe that SADA improves the results of all the considered baselines for both scenarios. More specifically, SADA attains a performance boost of up to mAP in the first scenario and in the second. Moreover, our method improves by mAP and mAP the performance of the second best-performing model in both scenarios, respectively. Interestingly, in both cases, the second best-performing method is DANN[1] which considerably improves other more complex methods like SSTDA[7]. Another interesting note is that WDGRL[42] presents a considerable variability in the performance. While this is the third best-performing method in the first scenario, its performance drops drastically in the second attaining only a mAP. This contrasts with the performance of similar methods like ADDA[40] or DANN[1].
@@ D.1 Domain adaptation methods
| MSTN [2] | 30.17 | 27.32 | 18.55 | 25.34 |
| SSTDA [7] | 29.58 | 26.81 | 18.36 | 24.92 |
| TranSVAE [65] | 28.31 | 25.67 | 16.72 | 23.57 |
@@
D.2 Ablation over variants of the SADA loss
@@
Closely related to Sec. IV.C of the main text, in Table S6 we ablate over different variants of our loss on scenario S2 and S4. These scenarios define the source domain as white old and new counter kitchens and the rest as the target, respectively. Observe in this table that our main insights from Sec. IV.C of the main text are mostly confirmed. Concretely, one of our main observations was that aligning the complete global distribution was generally preferable over aligning the partial background distribution only. This observation is satisfied in most of the scenarios that we study –with the exception of scenario S3. Similarly, in Sec. IV.C of the main text we argued that combining alignment loss that presented overlaps yielded also a decrease in performance. More in detail, this happens when we couple either the local or the background loss with a global alignment approach. In these cases, the embeddings have a double (simultaneous) alignment objective, which we argue is not desirable. As observed in Table S6, this observation is also mostly satisfied in these scenarios as the combination of these losses yields a performance degradation in 7 of the 8 cases studied in this paper.